ssume that the volume of a mitochondrion is um^3 and that the fraction of this volume occupied by the intermembrane space is 20%. The remaining 80% of the pace is a matrix (approximately). The pH within the matrix is found to be 7.8, and the pH within the intermembrane space is found to be 6.4.

a) how many free protons are expected to be in the intermembrane space?

b) how many free protons are expected to be in the matrix?

HINT be careful with unit conversion!
Answer #1

PH = - log (H+)

-pH = log (H+)

10 -pH = (H+)

10 -7.8 = 1.58 x 10 -8 M (a) free protons that are expected to be in the intermembrane space

10 -6.4 = 3.98 x 10 -7 M b) free protons that are expected to be in the matrix